QuinnLoves:

The graph below shows the cost of pens based on the number of pens in a pack. What would be the cost of one pen?

QuinnLoves:

1 attachment
jakfishman:

ok so what is the closest you have to 1 pen?

QuinnLoves:

3?

jakfishman:

yes so how would you find from 3 pens costing $4.50 to 1pen equals x?

QuinnLoves:

5.00?

QuinnLoves:

13.00?*

jakfishman:

no thats adding $0.50

QuinnLoves:

13.50?

jakfishman:

no you have to go down what do you have to do subtract or divide

QuinnLoves:

Subtract..?

jakfishman:

no divied you dived 4.5/3=x

QuinnLoves:

darn

jakfishman:

x/y=how much the pen cost

QuinnLoves:

4.5/3=1.5

QuinnLoves:

I belive

jakfishman:

yes

QuinnLoves:

Yess

QuinnLoves:

Thank you!

jakfishman:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @QuinnLoves Thank you! \(\color{#0cbb34}{\text{End of Quote}}\) no thank you for wanting to learn

QuinnLoves:

:)